OpenStudy (anonymous):

the radius of a circular puddle is growing at a rate of 20 cm/s. How fast is the area growing at the instant when it equals 25 cm^2? (use the area formula to determine the radius at that instant)

OpenStudy (anonymous):

\[ A=\pi r^2\\ \frac{dA}{dt} = 2\pi r\frac{dr}{dt} \] the radius of a circular puddle is growing at a rate of 20 cm/s\[ \frac{dr}{dt} = 20 \] the area growing at the instant when it equals 25 cm^2?\[ \pi r^2 = 25 \]

OpenStudy (agent0smith):

^ yep, now just find r (from that last equation) and put it into dA/dt.